Please help me!
I need to graph this and I have no idea how to do it.
x^2 -x-5 = 0 I need at least 4 points,
Thank you so much!

Answer by lwsshak3(11628)

I need to graph this and I have no idea how to do it.
x^2 -x-5 = 0 I need at least 4 points,
**
Like all 2nd degree equations, curve will be that of a parabola.
Rewrite equation:
y=x^2-x-5
y-intercept:
set x=0, then solve for y
y=-5
..
complete the square
y=(x^2-x+.25)-5-.25
y=(x-.5)^2-5.25
This is an equation for a parabola of the standard form: y=(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex.
vertex(0.5,-5.25)
..
x-intercepts
set y=0
(x-.5)^2-5.25=0
(x-.5)^2=5.25
take sqrt of both sides
x-.5=±√(5.25)≈±2.3
x≈.5±2.3
x≈2.8
or
x≈-1.8
You now have the following 4 points with which you can graph the given function:
(0,-5) y-intercept
(0.5,-5.25) vertex
(2.8,0) x-intercept
(-1.8,0) x-intercept